### v1

# Scalable Multi-view Clustering via Explicit Kernel Features Maps

Abstract
A growing awareness of multi-view learning as an important component in data science and machine learning is a consequence of the increasing prevalence of multiple views in real-world applications, especially in the context of networks. In this paper we introduce a new scalability framework for multi-view subspace clustering. An efficient optimization strategy is proposed, leveraging kernel feature maps to reduce the computational burden while maintaining good clustering performance. The scalability of the algorithm means that it can be applied to large-scale datasets, including those with millions of data points, using a standard machine, in a few minutes. We conduct extensive experiments on real-world benchmark networks of various sizes in order to evaluate the performance of our algorithm against state-of-the-art multi-view subspace clustering methods and attributed-network multi-view approaches.
1 Introduction
The explosive growth of data from diverse sources, such as social media, sensor networks, and online platforms, has led to the emergence of complex high-dimensional datasets. These datasets often contain multiple views of the same underlying data, each capturing different aspects or perspectives. Traditional clustering algorithms, designed for single-view data, face significant challenges in effectively capturing the intricate relationships and structures extending across multiple views. In this context, multi-view clustering, which integrates information from multiple views in order to enhance clustering performance, has emerged as a powerful paradigm. It has been successfully applied to various applications including computer vision [28, 25], natural language processing [12, 16], and bioinformatics [20, 23, 21]. For a more in-depth discussion, please refer to [1].
In this work we are specifically interested in multi-view subspace clustering [7, 22, 2]. By exploring the latent subspaces within each view and leveraging the complementary nature of different views, multi-view subspace clustering algorithms can reveal hidden structures that may not be apparent when considering individual views in isolation. This not only leads to more accurate clustering results, but also enables a deeper understanding of the underlying data. Specifically, multi-view subspace clustering for attributed network data comprising interconnected entities with associated attributes offers a powerful framework for analyzing complex datasets, and for this reason it has gained significant attention in recent literature [4, 13, 14, 17]. One prominent application of multi-view clustering for attributed network data is in social network analysis, given that social networks often exhibit rich attributes relating to individuals or groups – demographic information, interests, affiliations, and so on [4, 9]. Another application is in biological research, where multi-omics data can be integrated using multiple graph convolutional networks [24]. Finally, recommender systems often rely on multiple sources of information, such as user preferences, item attributes, and social connections [11, 27, 19].
The computational complexity of multi-view subspace clustering has hindered its widespread adoption in real-world applications, even though some efficient approaches have been proposed. These can involve, for example, anchor-based techniques for reducing the size of matrices used in optimization [10, 22]. However, tests that we performed show that existing methods often fail to achieve the desired level of efficiency and clustering performance in the case of large-scale multi-view datasets, and consequently that further advances in this area need to be made to unlock the full potential of multi-view subspace clustering.
In this paper, rather than using anchor-based methods, we tackle the scalable multi-view subspace clustering problem via a novel algorithm that is able to improve on the clustering accuracy and computational efficiency of existing approaches. Our algorithm leverages low-rank subspace clustering and properties of kernel feature maps in order to capture complex relationships between views and improve the robustness of clustering results. Similarly, to some state-of-the-art models, our approach scales linearly with the number of inputs, but it scales more favorably with other factors such as the number of dimensions and clusters. Additionally, it harnesses the inherent parallelism present in the multi-view setting such that large-scale datasets can be processed efficiently. In summary, the key contributions of this paper are threefold:
  1. 1.
    First, we introduce a novel framework that integrates multiple views of the data subspace structure graph into a unified consensus subspace structure graph, effectively capturing the complementary information across views and allowing for a new scalability framework for multi-view subspace clustering.
  2. 2.
    Second, we propose an efficient optimization strategy that scales well with very large-scale datasets. By leveraging the properties of kernel features maps, our algorithm significantly reduces the computational burden while maintaining state-of-the-art clustering performance. This scalability makes our approach suitable for handling real-world applications, since it is able to scale to datasets with millions of data points.
  3. 3.
    Third, we conduct extensive experiments, including statistical significance testing, on real-world benchmark networks of different scales and degrees of overlap, to evaluate our algorithm against state-of-the-art multi-view subspace clustering methods and attributed-network-specific multi-view approaches. Our experiments are entirely reproducible, ensuring transparency and enabling others to validate our results.
The remainder of this paper is structured as follows. In section 2 we discuss related works. Section 3 introduces the necessary background to develop our approach. In section 4 we develop our proposed method and explore the algorithm and its complexity. Section 5 contains a detailed description of our experiments. Finally, we give our conclusion in section 6. A schematic representation of our approach is shown in Figure 1.
Figure 1: Schematic diagram: the blue section represents the operations that are explicitly performed in our approach; the green section represents section represents what is transpiring implicitly (operations in the same column represent the same step). Given a set of inputs, we apply an SVD to obtain the left singular values (equivalent to implicitly computing a coefficient matrix using low-rank subspace clustering). We then apply a nonnegative feature map (equivalent to implicitly computing an affinity matrix). We then weigh and concatenate the obtained embeddings of each view (equivalent to summing the affinity matrices into a single consensus one). Finally, we perform an SVD followed by a k-Means to obtain a consensus partition (equivalent to performing spectral clustering on the consensus affinity matrix)
2 Related Work
This section discusses prior works relating to scalable multi-view subspace clustering. We also look at graph multi-view clustering, since we use attributed networks to evaluate performance.
2.1 Scalable Multi-view Subspace Clustering
In [14] Large-Scale Multi-View Subspace Clustering in Linear Time (LMVSC) [10] was introduced. This method, where smaller graphs are created for each view and then integrated in order to enable spectral clustering on a reduced graph size, derives from the concept of anchor graphs. The Scalable Multi-view Subspace Clustering with Unified Anchors (SMVSC) [22] approach integrates anchor learning and graph construction into a unified optimization framework, resulting in a more accurate representation of latent data distribution by learned anchors and a more discriminative clustering structure. In [2], an efficient orthogonal multi-view subspace clustering (OMSC) approach was proposed, featuring joint anchor learning, graph construction and partitioning. Recently, [29] developed a Flexible and Diverse Anchor Graph Fusion for Scalable Multi-view Clustering (FDAGF) that introduces a fusion strategy for multi-size anchor graphs.
2.2 Graph Multi-view Clustering
In [17], a multi-view contrastive graph clustering (MCGC) method was proposed, designed to learn a consensus graph with contrastive regularization over the learned graph matrix. [14] introduced a novel multi-view attributed graph clustering (MAGC) framework, which incorporates node attributes and graphs. It is an approach that stands out in three important respects: utilizing graph filtering for smooth node representation, rather than instead deep neural networks; learning a consensus graph from data to address noise and incompleteness in the original graph; and exploring high-order relations through a flexible regularizer design. The MvAGC model [13] is a simple yet effective approach for Multi-view Attributed Graph Clustering. It applies a graph filter to smooth the features without relying on neural network parameter learning. A reduced computational complexity is achieved through a novel anchor point selection strategy. In addition, a new regularizer is introduced to explore high-order neighborhood information. [6] introduced (LMGEC), a generic linear model for performing multi-view attributed graph representation learning and clustering simultaneously. It uses a simple weighing scheme based on inertia and a single-hop neighborhood propagation.
3 Preliminaries
In what follows, matrices are denoted in boldface uppercase letters and vectors in boldface lowercase letters. $[n]$ corresponds to the integer set $\{1,\ldots,n\}$. Tr is the trace operator. $\oplus$ denotes column-wise concatenation.
3.1 Subspace Clustering
Given a data matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$ and a desired number of clusters $k$, most subspace clustering problems have the following form:
$\min_{\mathbf{C}}\quad$ $\|\mathbf{X}-\mathbf{C}\mathbf{X}\|^{2}\,+\,\Omega(\mathbf{C})\quad\text{s.t.}\quad\mathbf{C}\in\mathcal{C}$ (1)
where $\mathbf{C}$ is called a coefficient matrix, $\Omega$ is a regularization function and $\mathcal{C}$ is the set of feasible solutions. Once a solution $\mathbf{C}$ is found, a subspace affinity graph represented using its adjacency matrix $\mathbf{W}$ is constructed, typically as $\mathbf{W}=(|\mathbf{C}|+|\mathbf{C}^{\top}|)/2$. Finally, a partition of $k$ groups is obtained via spectral clustering on $\mathbf{W}$. Different values of $\Omega$ and $\mathcal{C}$ can call for different subspace clustering techniques: these might, for example, be low-rank approaches [15] or sparse approaches [3].
$\min_{\mathbf{C}\in\mathcal{C}}\quad$ $\|\mathbf{X}-\mathbf{C}\mathbf{X}\|^{2}\,+\,\|\mathbf{C}\|_{*}\quad\text{s.t.}\quad\mathbf{C}=\mathbf{C}^{\top}$ (2)
Note that all norms here refer to the Frobenius norm, unless otherwise specified.
3.2 Scalable Subspace Clustering
In [5], the authors proposed a scalable low-rank subspace clustering model. Specifically, the problem they considered is:
$\min_{\mathbf{U}}\quad$ $\|\mathbf{X}-\mathbf{U}\mathbf{U}^{\top}\mathbf{X}\|^{2}\quad\text{s.t.}\quad\mathbf{U}^{\top}\mathbf{U}=\mathbf{I}.$ (3)
To address the problem they use nonnegative kernel maps to create an affinity matrix $\mathbf{W}=\Phi(\mathbf{U})\Phi(\mathbf{U})^{\top}.$ The advantage of computing the affinity matrix in this manner is that spectral partitioning can be obtained via singular values decomposition on $\Phi(\mathbf{U})$ instead of performing eigendecomposition on $\mathbf{W}$, leading to substantial gains in spatial and computational efficiencies.
3.3 Multi-view Subspace Clustering
Given a set of features $\mathbf{X}_{1},\ldots,\mathbf{X}_{V}$ describing the same data points, a rudimentary multi-view subspace clustering model may consider a problem of the following form:
$\min_{\mathbf{F},\mathbf{C}_{1},\ldots,\mathbf{C}_{V}}\quad$ $\sum_{v}\|\mathbf{X}_{v}-\mathbf{C}_{v}\mathbf{X}_{v}\|^{2}+\texttt{Tr}(\mathbf{F}^{\top}\mathbf{C}_{v}\mathbf{F})$ (4)
such that $\mathbf{F}\in\{0,1\}^{n\times k},\quad\mathbf{F}\mathbf{1}=\mathbf{1}$
$\mathbf{C}_{1},\ldots,\mathbf{C}_{V}\in\mathcal{C},$
where $\mathbf{C}_{v}$ is the coefficient matrix of the $v$-th view and $\mathbf{F}$ is a a consensus partition matrix. However, although the block structures in different view-specific coefficient matrices are similar, the magnitudes and the signs of the entries in each $C_{v}$ can vary greatly. To solve this issue, [7] proposed using the affinity matrices $\mathbf{W}_{v}$ rather than the coefficient matrices $\mathbf{C}_{v}$. We generalize their formulation by dropping the specific constraints that the authors used. This results in the following generic multi-view subspace clustering problem:
$\min_{\mathbf{F},\mathbf{C}_{1},..,\mathbf{C}_{V}}$ $\quad\sum_{v}\|\mathbf{X}_{v}-\mathbf{C}_{v}\mathbf{X}_{v}\|^{2}\quad+\texttt{Tr}\left(\mathbf{F}^{\top}\left(\mathbf{I}-\tfrac{|\mathbf{C}_{v}|+|\mathbf{C}_{v}^{\top}|}{2}\right)\mathbf{F}\right)$ (5)
such that $\mathbf{F}\in\{0,1\}^{n\times k},\quad\mathbf{F}\mathbf{1}=\mathbf{1},\mathbf{C}_{1},\ldots,\mathbf{C}_{V}\in\mathcal{C}.$
Once solutions $\mathbf{C}_{1},\ldots,\mathbf{C}_{V}$ are available, depending on the regularization used, spectral clustering can be performed on $\mathbf{W}=\sum_{v}\mathbf{C}_{v}$ to obtain the partition matrix $\mathbf{F}$.
4 Proposed Methodology
We consider the multi-view framework presented in problem 5 and like in 3 we introduce into it low-rank constraints, also using the same affinity matrix. However, here we are considering a symmetrically normalized affinity matrix rather than the unnormalized one. We obtain the following problem:
$\min_{\mathbf{F},\mathbf{U}_{1},\ldots,\mathbf{U}_{V}}\quad$ $\sum_{v}\|\mathbf{X}_{v}-\mathbf{U}_{v}\mathbf{U}_{v}^{\top}\mathbf{X}_{v}\|^{2}+\ \texttt{Tr}(\mathbf{F}^{\top}(\mathbf{I}-\mathbf{W}_{v})\mathbf{F})$ (6)
such that $\mathbf{W}_{v}=\mathbf{D}_{v}^{-\frac{1}{2}}\Phi(\mathbf{U}_{v})\Phi(\mathbf{U}_{v})^{\top}\mathbf{D}_{v}^{-\frac{1}{2}}$
$\mathbf{F}\in\{0,1\}^{n\times k},\quad\mathbf{F}\mathbf{1}=\mathbf{1}$
where $\mathbf{U}\in\mathbb{R}^{n\times f}$ is a low-rank matrix, and $\mathbf{W}_{v}$ is the normalized affinity matrix of view $v$. If we were to solve this problem directly, the consensus partition $\mathbf{F}$ would be obtained via spectral clustering on
$\mathbf{W}=\sum_{v}\mathbf{W}_{v}=\sum_{v}\mathbf{D}_{v}^{-\frac{1}{2}}\Phi(\mathbf{U}_{v})\Phi(\mathbf{U}_{v})^{\top}\mathbf{D}_{v}^{-\frac{1}{2}}.$ (7)
The problem of applying spectral clustering via the proper decomposition of this matrix would result in a computational complexity of $\mathcal{O}(n^{2}k)$ and a spatial complexity of $\mathcal{O}(n^{2})$. This means that we would lose the efficiency argued by [5] and thus lose the advantage of using such an approach in the first place. Therefore, we propose solving this problem via the summation property of kernels.
4.1 Weighing Views with respect to their Clustering Structure
We additionally introduce a weighting scheme on the view-specific terms in problem 6, since the different views should not always have the same contribution level towards the consensus affinity matrix. Here, we evaluate the importance of each view using the affinity matrix corresponding to that view. Specifically, we consider the following problem:
$\min_{\mathbf{F},\mathbf{U}_{1},\ldots,\mathbf{U}_{V}}$ $\sum_{v}\|\mathbf{X}_{v}-\mathbf{U}_{v}\mathbf{U}_{v}^{\top}\mathbf{X}_{v}\|^{2}\ +\ \lambda_{v}\ \ \texttt{Tr}(\mathbf{F}^{\top}(\mathbf{I}-\mathbf{W}_{v})\mathbf{F})$ (8)
such that $\mathbf{W}_{v}=\mathbf{D}_{v}^{-\frac{1}{2}}\Phi(\mathbf{U}_{v})\Phi(\mathbf{U}_{v})^{\top}\mathbf{D}_{v}^{-\frac{1}{2}}$
$\mathbf{F}\in\{0,1\}^{n\times k},\quad\mathbf{F}\mathbf{1}=\mathbf{1}$
where we define the regularization vector $\boldsymbol{\lambda}=[\lambda_{1},\ldots,\lambda_{V}]$ with respect to the cluster-ability of the $v$-th view:
$\qquad\lambda_{v}=\frac{\texttt{exp}(\ \texttt{Tr}(\ \mathbf{G}_{v}^{\top}(\mathbf{I}-\mathbf{W}_{v})\mathbf{G}_{v})\ /\ T\ )}{\sum_{w}\texttt{exp}(\ \texttt{Tr}(\ \mathbf{G}_{w}^{\top}(\mathbf{I}-\mathbf{W}_{w})\mathbf{G}_{w})\ /\ T\ )}$ (9)
$\text{such that}\quad\forall w\in[V]\qquad\ \ \mathbf{W}_{v}=\Phi(\mathbf{U}_{w})\Phi(\mathbf{U}_{w})^{\top}$
$\qquad\qquad\qquad\qquad\qquad\quad\mathbf{G}_{w}\in\{0,1\}^{n\times k},\quad\mathbf{G}_{w}\mathbf{1}=\mathbf{1}$
where $T$ is a temperature parameter and $\mathbf{G}_{1},\ldots,\mathbf{G}_{V}$ are different view-specific partition matrices that for spectral clustering we solve by applying singular value decomposition on $\mathbf{D}_{w}^{-\frac{1}{2}}\Phi(\mathbf{U}_{w})$ and then applying a clustering algorithm on the resulting left singular vectors and repeating the process for each $v\in[V]$.
4.2 Multi-view Scalability via Kernel Summation
Since we are using kernels to compute the view-specific affinity matrices $\mathbf{C}_{1},\ldots,\mathbf{C}_{V}$, we propose using another property of kernels in order to obtain a factorized matrix consensus affinity matrix, which will then allow for efficient spectral clustering. Consider the following:
Property 1.
Given two kernels $k_{a}$ and $k_{b}$ then their summation $k(x,y)=k_{a}(x,y)+k_{b}(x,y)$ is also a kernel.
This means that matrix 7 is also a kernel and, consequently, has a decomposition of the form $\mathbf{B}\mathbf{B}^{\top}$ for some matrix $\mathbf{B}$ (this is property of kernel matrices). Additionally, we have the obvious fact:
Property 2.
When $k_{a}$ and $k_{b}$ are nonnegative kernels, i.e. $\forall x,y$ $k_{a}(x,y)\geq 0$ and $k_{b}(x,y)\geq 0$, then $k(x,y)=k_{a}(x,y)+k_{b}(x,y)$ is also a nonnegative kernel.
This means that every entry of 7 is nonnegative, making it a valid affinity matrix. To solve the efficiency issue of combining the different view-specific affinity matrices we therefore need a way of obtaining a vector realization $\mathbf{B}$ for the decomposition of 7. Once we have $\mathbf{B}$, we can obtain a spectral clustering via singular value decomposition of $\mathbf{B}$ instead of eigendecomposition of $\mathbf{B}\mathbf{B}^{\top}$. To this end we make use of the following property:
Property 3.
Given kernels $k_{a}$ and $k_{b}$ and their associated feature maps $\Phi_{a}$ and $\Phi_{b}$, the feature map corresponding to their summation kernel $k$ is $\Phi(x)=\Phi_{a}(x)\oplus\Phi_{b}(x)$.
Using this property, we can infer that a vector realization is $\mathbf{B}=\left[\mathbf{D}_{1}^{-\frac{1}{2}}\Phi(\mathbf{U}_{1}),\ldots,\mathbf{D}_{V}^{-\frac{1}{2}}\Phi(\mathbf{U}_{V})\right]$. This decomposition means that the complexity of the approach is proportional to dimension $m$ of the feature map $\Phi$. This leads to large speedup and memory saving when $m<<n$.
4.3 Optimization
In addition to the use of properties associated with kernel feature maps, the efficiency of our approach also stems from the fact that we do not use an iterative algorithm, in contrast to many other state-of-the-art models, e.g, [17]. The algorithm consists of the following steps:
4.3.1 Solving for $\mathbf{C}_{1},\ldots,\mathbf{C}_{V}$
The first step is fixing
$\mathbf{C}_{1},\ldots,\mathbf{C}_{v-1},\mathbf{C}_{v+1},\ldots,\mathbf{C}_{V}$ and $F$ in problem 6 and reiterating for each $v\in[V]$. The resulting problem is 3, and the solution is to set $\mathbf{U}$ as the $f$ left singular vectors of $\mathbf{X}$ associated with the largest singular values. Note that to solve our approach in one step and for scalability, as opposed to [7], we consider $\mathbf{W}_{v}$ not to be a function of $\mathbf{C}_{v}$.
4.3.2 Computing $\lambda_{1},\ldots,\lambda_{V}$
Once we have obtained the singular vectors associated with each coefficient matrix $\mathbf{C}_{v}$, we compute the corresponding $\lambda_{v}$ using Formula 9.
4.3.3 Solving for $\mathbf{F}$
When fixing every $\mathbf{C}_{v}$, we have to solve the following problem
$\min_{\mathbf{F}}\quad$ $\sum_{v}\texttt{Tr}(\mathbf{F}^{\top}(\mathbf{I}-\mathbf{W}_{v})\mathbf{F})$ (10)
such that $\mathbf{F}\in\{0,1\}^{n\times k},\quad\mathbf{F}\mathbf{1}=\mathbf{1}.$
The solution, then, consists in using property 3, and noticing that:
$\mathbf{W}=$ $\sum_{v}\mathbf{W}_{v}=\sum_{v}\lambda_{v}\mathbf{D}_{v}^{-\frac{1}{2}}\Phi(\mathbf{U}_{v})\Phi(\mathbf{U}_{v})^{\top}\mathbf{D}_{v}^{-\frac{1}{2}}$ (11)
$=$ $\left(\oplus_{v}\ \sqrt{\lambda_{v}}\mathbf{D}_{v}^{-\frac{1}{2}}\Phi(\mathbf{U}_{v})\right)\left(\oplus_{v}\ \sqrt{\lambda_{v}}\mathbf{D}_{v}^{-\frac{1}{2}}\Phi(\mathbf{U}_{v})\right)^{\top}.$
The solution $\mathbf{F}$ of this problem is obtained by applying $k$-means clustering on the left singular values corresponding to the largest $k$ singular values of matrix $\oplus_{v}\ \sqrt{\lambda_{v}}\mathbf{D}_{v}^{-\frac{1}{2}}\Phi(\mathbf{U}_{v})$.
A pseudo-code for our approach is given in algorithm 1.
4.4 Complexity Analysis
To complement our intuitive justification for the efficiency of our approach, we give a theoretical analysis of its complexity:
  1. 1.
    Solving for $\mathbf{C}_{1},\ldots,\mathbf{C}_{V}$ This process takes $O(vnd\log(f))$ where $f$ is the chosen rank due to the randomized singular values decomposition.
  2. 2.
    Computing $\lambda_{1},\ldots,\lambda_{V}$ The complexity corresponds to computing the projections using feature map $\Phi$. We consider the case of using a quadratic polynomial feature map whose computational complexity is $O(nf^{2})$. The computational complexity of this step is then $O(Vnf^{2}log(k))$.
  3. 3.
    Solving for $\mathbf{F}$ Solving for $\mathbf{F}$ Finally, the process of computing the concatenated matrix $\mathbf{B}$ and then applying singular value decomposition is $O(Vnf^{2}\log(k))$. Applying k-means afterwards results in $O(tVnf^{2}k)$ operations.
  4. 4.
    The effect of the choice of the non-negative Kernel It is possible to use nonnegative kernels other than the quadratic polynomial. However, in the case of infinite-dimensional kernels such as RBF and sigmoid kernels or, in the case of polynomials, kernels with large even degrees, it is necessary to use kernel feature map approximation techniques such as Nystroem [26] or polynomial sketching [18]. In these cases, the factor $f^{2}$ is replaced by the dimension of the feature map approximation, and the complexity of computing the approximation of the kernel has to be added to the total complexity of our approach.
In table 1 we report the complexity of several state-of-the-art algorithms. We see that considerably more factors must be taken into consideration than are needed when using our approach. Note that we have added the complexity of learning graph representations (using 12) to non-graph methods.
Data: Data matrices $\mathbf{X}_{1},\ldots,\mathbf{X}_{V}$, number of clusters $k$, number of components $f$, kernel feature map $\Phi$, temperature $T$
Result: Consensus partition matrix $\mathbf{F}$
for  $v\in[V]$ do
       Subtract column means from $\mathbf{X}_{v}$;
       Construct matrix $\mathbf{U}_{v}$ by stacking the $f$ first left singular vectors of $\mathbf{X}_{v}$ as its columns;
       Compute $d$ the row sums of $\mathbf{B}_{v}\mathbf{B}_{v}^{\top}$ efficiently;
       $\mathbf{B}_{v}=\mathbf{B}_{v}*\textbf{d}^{-0.5}$;
       Construct matrix $\mathbf{Q}_{v}$ by stacking the $f+1$ first left singular vectors of $\mathbf{B}_{v}$ as its columns and then drop the first column;
       Compute partition matrix $\mathbf{G}_{v}$ via k-means on $\mathbf{Q}_{v}$ with $k$ clusters
Compute $\lambda_{1},...,\lambda_{V}$ using Formula 9 with temperature $T$, $\mathbf{G}_{v}$ and $\mathbf{W}_{v}=\mathbf{B}_{v}\mathbf{B}_{v}^{\top}$ ;
$\mathbf{B}=\oplus_{v}\lambda_{v}\mathbf{B}_{v}$ ;
Compute $d$ the row sums of $\mathbf{B}\mathbf{B}^{\top}$ efficiently ;
$\mathbf{B}=\mathbf{B}*\textbf{d}^{-0.5}$ ;
Construct matrix $\mathbf{Q}\phantom{v}$ by stacking the $f+1$ first left singular vectors of $\mathbf{B}\phantom{v}$ as its columns and then drop the first column ;
Compute partition matrix $\mathbf{F}$ via k-means on $\mathbf{Q}$ with $k$ clusters
Algorithm 1 MvSCK
Table 1: Complexity of some of the considered approaches on network data. $m$ represents the number of anchors. $|E|$ is the number of edges in the graphs used in Formula 12. For the sake of simplicity, we consider that the dimension of the features and the number of graph edges in the different views are the same.
Method Time Complexity
LMVSC $\mathcal{O}(Vp|E|d+nm^{3}V+m^{3}V^{3}+mVn+nk^{2}t)$
FDAGF $\mathcal{O}(Vp|E|d+rVdm^{2}+rm^{2}n+rVdmn)$
SMVSC $\mathcal{O}(Vp|E|d+Vd^{3}+Vd^{2}k^{2}+md^{2}+dmk^{2}+nm^{3})$
MAGC $\mathcal{O}(Vn^{2}d)$
MvAGC $\mathcal{O}(tm^{3}V+tVm^{2}nd)$
LMGEC $\mathcal{O}(V|E|d+Vndk)$
MvSCK $O(Vp|E|d+Vnf^{2}\log(k))$
Table 2: Characteristics of the Datasets. The imbalance is given as the ratio between the cardinalities of the most frequent and least frequent classes.
Dataset Scale Nodes Features Graph and Edges Clusters Imbalance
ACM Small 3,025 1,830
Co-Subject (29,281)
Co-Author (2,210,761)
3 1.1
DBLP Small 4,057 334
Co-Author (11,113)
Co-Conference (5,000,495)
Co-Term (6,776,335)
4 1.6
Amazon Photos Medium 7,487
745
7,487
Co-Purchase (119,043) 8 5.7
Amazon Computers Medium 13,381
767
13,381
Co-Purchase (504,937) 10 17.5
OGBN-ArXiv Large 169,343 128
k-NN (1,862,773)
Co-Citation (2,484,941)
40 942.1
OGBN-Products Large 2,449,029 100
k-NN (26,939,319)
Co-Purchase (126,167,053)
47 668,950
Table 3: Clustering results on the small scale datasets.
ACM DBLP
CA CF1 NMI ARI CA CF1 NMI ARI
LMVSC 90.33 ±0.0 90.59 ±0.1 68.33 ±0.2 73.64 ±0.1 52.72 ±0.0 58.30 ±0.0 20.15 ±0.0 14.97 ±0.0
OMSC 70.21 ±0.0 71.11 ±0.0 45.81 ±0.0 43.55 ±0.0 91.45 ±0.0 90.95 ±0.0 74.59 ±0.0 79.82 ±0.0
FDAGF 70.27 ±0.0 79.39 ±0.0 47.07 ±0.1 45.20 ±0.1 89.61 ±0.0 89.39 ±0.0 71.56 ±0.0 75.78 ±0.0
SMVSC 70.45 ±0.0 71.80 ±0.0 46.35 ±0.0 43.91 ±0.0 91.45 ±0.0 90.87 ±0.0 74.50 ±0.0 79.88 ±0.0
MvAGC 83.35 ±4.9 83.49 ±5.0 55.86 ±7.1 58.78 ±8.6 76.62 ±6.4 67.10 ±3.5 59.99 ±10.5 59.87 ±6.5
MAGC 55.81 ±7.8 47.58 ±4.1 26.09 ±12.6 20.05 ±16.6 91.93 ±0.6 91.43 ±0.6 75.28 ±1.1 80.53 ±1.4
MCGC 89.92 ±0.0 89.93 ±0.0 67.03 ±0.0 72.56 ±0.0 87.63 ±0.0 77.24 ±9.9 69.37 ±2.0 79.24 ±7.9
LMGEC 93.00 ±0.0 93.06 ±0.0 75.13 ±0.1 80.27 ±0.1 92.85 ±0.0 92.37 ±0.0 77.40 ±0.0 82.84 ±0.0
MvSCK 93.21 ±0.1 93.22 ±0.1 74.97 ±0.1 80.76 ±0.1 93.09 ±0.1 92.57 ±0.1 78.05 ±0.2 83.36 ±0.2
Table 4: Clustering results on the medium scale datasets.
Amazon Photos Amazon Computers
CA CF1 NMI ARI CA CF1 NMI ARI
LMVSC 61.23 ±1.2 55.87 ±1.4 57.17 ±0.7 36.60 ±0.7 64.97 ±0.9 62.24 ±3.4 52.39 ±0.9 48.76 ±0.1
OMSC 61.44 ±0.0 N/A 49.28 ±0.0 36.73 ±0.0 63.99 ±0.0 N/A 42.27 ±0.0 42.89 ±0.0
FDAGF 55.48 ±0.1 49.37 ±0.0 52.96 ±0.1 30.01 ±0.1 64.68 ±4.1 56.93 ±0.8 52.82 ±0.4 44.94 ±2.6
SMVSC 61.67 ±0.1 55.74 ±0.3 50.83 ±0.2 35.99 ±0.1 64.72 ±0.1 49.55 ±0.6 49.09 ±0.4 46.01 ±0.1
MvAGC 45.28 ±6.7 43.15 ±4.4 29.96 ±5.9 20.02 ±3.9 45.82 ±4.1 45.12 ±4.8 28.45 ±5.5 18.98 ±3.1
MAGC 52.36 ±2.7 43.76 ±3.5 48.86 ±2.5 21.42 ±2.3 63.48 ±6.1 51.81 ±10.2 52.47 ±10.1 35.48 ±9.4
MCGC 41.86 ±3.5 26.75 ±7.7 15.16 ±10.6 23.61 ±9.4 45.29 ±0.7 34.17 ±8.2 22.24 ±7.4 26.96 ±10.4
LMGEC 70.93 ±0.0 64.83 ±0.0 60.88 ±0.0 51.05 ±0.0 46.10 ±1.5 40.21 ±0.8 46.14 ±0.9 27.61 ±0.4
MvSCK 75.66 ±2.2 66.56 ±5.0 70.66 ±1.3 58.65 ±2.0 66.34 ±1.5 56.23 ±0.8 57.15 ±0.8 46.27 ±2.0
Table 5: Clustering results on the large scale datasets. We allow for a maximum runtime of two hours per
run otherwise we report a timeout.
OGBN-ArXiv OGBN-Products
CA CF1 NMI ARI CA CF1 NMI ARI
LMVSC 30.94 ±0.6 23.51 ±4.6 31.92 ±1.0 20.07 ±0.5 Timeout
OMSC Timeout Timeout
FDAGF Timeout Timeout
SMVSC Timeout Timeout
MvAGC Out of Memory Out of Memory
MAGC Out of Memory Out of Memory
MCGC Timeout Timeout
LMGEC 29.61 ±0.8 16.44 ±0.7 34.53 ±0.3 19.86 ±0.7 36.09 ±1.1 16.23 ±0.7 43.64 ±0.7 21.60 ±0.5
MvSCK 48.29 ±1.8 28.46 ±0.9 46.99 ±0.5 39.15 ±1.2 47.14 ±1.5 16.89 ±1.3 50.80 ±0.8 32.24 ±1.3
Table 6: Execution times on the different datasets (in seconds).
ACM DBLP Amazon Photos Amazon Computers OGBN-ArXiv OGBN-Products
LMVSC 2.36 ±0.1 2.97 ±0.1 37.78 ±3.7 144.26 ±14.2 2952.97 ±210.43 Timeout
OMSC 24.66 ±0.2 27.51 ±0.1 288.32 ±43.1 414.94 ±3.7 Timeout Timeout
FDAGF 37.31 ±1.3 68.28 ±3.2 382.76 ±4.2 1555.28 ±50.6 Timeout Timeout
SMVSC 16.94 ±1.6 29.68 ±0.6 169.55 ±14.5 352.40 ±95.5 Timeout Timeout
MvAGC 5.05 ±0.1 5.65 ±0.0 36.59 ±1.0 36.76 ±0.7 Out of Memory Out of Memory
MAGC 21.83 ±1.6 38.12 ±3.8 304.23 ±22.6 8699.82 ±16343.2 Out of Memory Out of Memory
MCGC 738.54 ±4.5 358.56 ±13.5 3803.44 ±434.1 11393.04 ±1690.1 Timeout Timeout
LMGEC 0.63 ±0.4 0.51 ±0.5 2.95 ±0.3 7.80 ±1.0 44.87 ±1.4 601.32 ±29.2
MvSCK 0.18 ±0.1 0.19 ±0.1 0.86 ±0.1 5.96 ±0.2 20.43 ±3.8 239.32 ±14.2
Table 7: MvSCK results and running times (in seconds) using different kernels.
ACM DBLP Photos Computers ArXiv Products
Kernel CA Time CA Time CA Time CA Time CA Time CA Time
Quadratic 93.2 0.2 93.1 0.2 75.7 0.9 66.3 6.0 48.3 20.4 47.1 239.3
RBF 93.2 0.8 93.1 2.0 76.2 2.4 65.0 6.5 48.7 15.8 18.6 325.2
Sigmoid 93.2 0.9 93.0 1.9 74.7 2.3 65.6 6.4 34.0 16.6 33.2 231.3
(a) ACM
(b) DBLP
(c) Amazon Photos
(d) Amazon Computers
(e) OGBN-ArXiv
(f) OGBN-Products
Figure 2: Sensitivity of MvSCK w.r.t the temperature parameter $T$. Note that in our experiments, we used $T=0.1$ for all datasets.
5 Experiments
In the following comparisons, our approach is identified with the label MvSCK (for Multi-view Subspace Clustering with explicit Kernel feature maps). We aim to answer the following research questions:
  • How does MvSCK compare with state-of-the-art in terms of:
    • RQ1: Clustering performance?
    • RQ2: Computational efficiency?
  • RQ3: Does our approach consistently outperform the state-of-the-art ?
  • RQ4: How do the hyper-paramters affect MvSCK?
  • RQ5: Is the view-importance parameter necessary?
5.1 Datasets
We evaluate our method on six benchmark datasets with different degrees of overlap: two small scale datasets, namely ACM and DBLP [4]; two medium datasets, namely Amazon Photos and Amazon Computers [19]; and two large datasets OGBN-ArXiv and OGBN-Products [9]. The six datasets are described briefly below:
  • ACM is a network dataset of research papers, where the associated node features are the papers’ keywords in bag-of-words format. The “Co-Subject” and “Co-Author” graphs respectively connect papers with the same theme or the same author.
  • DBLP is a network dataset of authors with bag-of-words node features representing keywords. “Co-Author”, “Co-Conference”, and “Co-Term” links indicate that the authors concerned have collaborated on papers, published papers at the same conference, or published papers with the same terms.
  • Amazon Photos & Amazon Computers are segments of the Amazon co-purchase network dataset. The nodes represent goods, and their attribute features are represented by bag-of-words of product reviews. The edges indicate that two goods are purchased together. The second set of attributes represents the word co-occurrence matrix.
  • OGBN-ArXiv represents a citation network encompassing a set of computer science arXiv papers. Each paper comes with a 128-dimensional feature vector obtained by averaging the embeddings of words in its title and abstract. We add a second $k$-nearest neighbors (k-NN) graph with 10 neighbors and added self-connections.
  • OGBN-Products is a graph representing an Amazon product co-purchase network. An edge between two products signifies that those products are purchased together. The node features are created by extracting bag-of-words features from the product descriptions and reducing their dimensionality to 100 via PCA. Similarly, to OGBN-ArXiv, we add a second k-NN graph.
The statistical summaries for the different datasets are given in Table 2.
5.2 Experimental Setup
Experimental setting
For the purposes of comparison, we validate our approach against the state-of-the-art approaches discussed in the Related Work section above, with reference to clustering accuracy (CA), clustering F1-score (CF1), adjusted rand index (ARI), and normalized mutual information (NMI). We use the official implementations of compared approaches. Parameters are set based on the recommendations of the authors. When these recommendations are not available, we set parameters similarly to how they were set on datasets of similar dimensions. To make comparisons fair we remove any supervised signals when running experiments in the original code; these would include, for example, the validation of runs based on supervised metrics such as accuracy or setting random seeds. For our approach, the kernel used is the quadratic kernel that was discussed in [5]. The number of components parameter $f$ was set equal to the number of clusters on all datasets, with the exception of OGBN-Products, where it was set to 10 due to its large size. Note that all metrics reported are averaged over five different runs. All experiments were performed on a laptop equipped with a 12th Gen Intel(R) Core(TM) i9-12950HX (24 CPU threads) 2.3 GHz processor, a 12GB GeForce RTX 3080 Ti GPU, and 64GB of RAM.
Learning Node Representations
To perform clustering on network datasets using generic approaches such as ours in addition to LMVSC, SMVSC, FDAGF and OMSC, we first need to learn node embeddings. Given a multi-view attributed network consisting of adjacency matrices $\mathbf{A}_{1},\ldots,\mathbf{A}_{V}$ and feature matrices $\mathbf{X}_{1},\ldots,\mathbf{X}_{V}$, we learn embeddings via Laplacian smoothing. This is a well-known technique that has been used in works such as [17, 6] to handle graph data. It involves consecutive neighborhood averaging steps:
$\forall v\in[V]\qquad\mathbf{X}_{v}\leftarrow\mathbf{A}_{v}^{p}\mathbf{X}_{v}$ (12)
where $p$ is the propagation order. Specifically, in our experiments, we set the propagation order as follows: 2 for ACM, 2 for DBLP, 20 for Amazon Photos, 40 for Amazon Computers, 60 for OGBN-ArXiv, and 100 for OGBN-Products. This step can be seen as a preprocessing step that is independent of the approaches themselves. It is a methodology that has emerged from work done on simplifying graph convolutional networks [27].
5.3 Clustering Results (RQ1)
The clustering outcomes for small, medium, and large graph datasets are detailed in Tables 3, 4, and 5 respectively. A comprehensive analysis of these tables reveals the superior performance of our approach relative to other methodologies across a majority of datasets and evaluation metrics. Remarkably, our method exhibits superiority even when compared to specialized techniques designed explicitly for graph datasets. Noteworthy observations include LMVSC’s competitiveness with our approach on the Amazon Computers dataset, where both methods achieve comparable results. Basically, while LMVSC outperforms our approach in terms of F1 and ARI metrics, our method outperforms it in accuracy and NMI metrics. This comparison underscores the strengths of our approach, showcasing its robust performance across diverse datasets and metrics. We observed some disparities between results reported in original papers and those obtained by running the code on the same datasets. Additionally, many approaches struggle to scale to challenging datasets like OGBN-ArXiv and OGBN-Products. Specifically, OMSC fails to retrieve the correct number of clusters in Amazon Photos and Computers, impacting its ability to compute CF1 scores.
5.4 Running Times (RQ2)
The detailed breakdown of running times for different models across all datasets is presented in Table 6, where our approach emerges as the frontrunner with markedly shorter execution times across small, medium, and large graph datasets. The notable efficiency of our method becomes especially evident when compared to alternative approaches. Notably, for the challenging OGBN-ArXiv and OGBN-Products datasets, our approach not only maintains its speed advantage but also adheres to the strict three-hour time limit we set. In contrast, several competing models exceed this limit, highlighting the computational efficiency of our proposed method in the face of diverse datasets.
5.5 Statistical Significance Testing (RQ3)
Figure 4 shows the results of the Holm mean-rank post-hoc test [8] over the clustering results. We compute the mean rank over rankings generated based on each (dataset, metric, run) triplet. We see that our approach is ranked first by a large margin, followed by LMGEC and LMVSC as second and third. After these three come various overlapping groups of approaches.
In Figure 5, we report the results of the Holm mean-rank post-hoc test over the execution times of the different algorithms. Ours is the most efficient, followed by LMGEC. LMVSC and MvAGC are ranked third, followed by SMVSC, OMSC, and then a group composed of MAGC and FDAGF. MCGC comes last.
Figure 3: Clustering accuracy with and without the regularization vector $\boldsymbol{\lambda}=[\lambda_{1},\ldots,\lambda_{V}]$.
5.6 Sensitivity Analysis (RQ4)
Temperature Parameter $\ T$
In our experiments we tried various values for the temperature hyperparameter. Figure 2 shows the performance of LMGEC for different temperature values on the different datasets and for the different clustering metrics. We see that on most datasets the performance remains mostly constant but that larger values deteriorate the performance. As such, we advise using $T\leq 1$. In our exeperiments, we set it to $T=0.1$ for all datasets.
Kernel Feature Map
In table 7 we report the results of our approach using different kernels, namely an exact quadratic kernel, Nystroem approximations of the radial basis function (RBF), and a sigmoid kernel. For the approximations we use $10*k$ components, apart from on the OGBN-Products dataset, which is too large and for which we use $100$ components only. The results show that using different kernels gives similar performances, except on OGBN-Products, where the dimension of the approximations would appear to be insufficient, leading to a poor approximation of the feature maps; the same is true for the sigmoid kernel on ArXiv. Note that in the context of RBF and sigmoid we use approximations, given that their exact feature maps are infinite-dimensional.
5.7 Ablation relative to the View-Importance Parameter $\boldsymbol{\lambda}$ (RQ5)
Figure 3 illustrates the impact of the regularization vector $\boldsymbol{\lambda}=[\lambda_{1},\ldots,\lambda_{V}]$ on various datasets. The influence of $\boldsymbol{\lambda}$ is evident in its ability to enhance performance across five out of the six datasets analyzed during ablation studies. Notably, there is no observed deterioration in performance on the remaining ACM dataset, indicating that the regularization vector does not adversely affect outcomes in this particular case.
11113333555577779999MvSCKLMGECLMVSCSMVSCFDAGFMAGCOMSCMCGCMvAGC
Figure 4: Holm post-hoc mean rank test ($\alpha=0.01$) with respect to clustering performance.
11113333555577779999MvSCKLMGECLMVSCMvAGCSMVSCOMSCMAGCFDAGFMCGC
Figure 5: Holm post-hoc mean rank test ($\alpha=0.01$) with respect to running times.
6 Conclusion
In conclusion, this paper introduced a novel scalable framework for multi-view subspace clustering, leveraging kernel feature maps in order to ensure an efficient computation of the consensus subspace affinity graph. In virtue of the properties intrinsic to kernel summation, the approach that we propose results in performance gains and significant speedup. Extensive experiments on real-world benchmark networks, whatever the degree of overlap, demonstrate the superiority of our algorithm over state-of-the-art methods, especially in the context of very large network datasets, where most other models used in our comparison failed to scale despite using the same computational resources.